OpenStudy (anonymous):

length of a 48 degree arc of a circle with a 12m radius

OpenStudy (anonymous):

48*12/360=1.6

OpenStudy (anonymous):

my options are 3.2\[\pi\] 4.25 pi 11.25pi and 12 pi

OpenStudy (anonymous):

1.6*2 pi=3.2 pi

OpenStudy (anonymous):

sorry for that...